OpenStudy (anonymous):

Which point lies on the graph of the piece wise function y = –x if x < 0 y = x2 + 1 if x 0? (–1, 2) (1, 0) (–2, 2) (0, –1)

OpenStudy (zzr0ck3r):

-2<0 and 2=-(-2)

OpenStudy (zzr0ck3r):

only 1 x > 0 and it does not work

OpenStudy (nubeer):

hmm should be the second option as have no condition given for x>1

OpenStudy (zzr0ck3r):

I think the end of his question is x>0

OpenStudy (zzr0ck3r):

or x>=0 either way (-2,2) works

OpenStudy (nubeer):

hmm i asked him he said it's x = 0

OpenStudy (zzr0ck3r):

then again (-2,2) is the only one that works

OpenStudy (anonymous):

Which point lies on the graph of the following piecewise function? y = –x if x < 0 y = x2 + 1 if x 0

OpenStudy (nubeer):

hmm true..

OpenStudy (anonymous):

Which point lies on the graph of the following piecewise function? y = –x if x < 0 y = x2 + 1 if x greater than or equal to 0

OpenStudy (anonymous):

thats the exact question

OpenStudy (zzr0ck3r):

only b) works

OpenStudy (nubeer):

well still the same answer as @zzr0ck3r mentioned.

OpenStudy (anonymous):

so 1,0?

OpenStudy (zzr0ck3r):

(-2,2)

OpenStudy (anonymous):

IT WAS RIGHT! Thanks a million.

OpenStudy (zzr0ck3r):

do you see what we did? y = –x if x < 0 y = x2 + 1 if x 0? (–1, 2) (1, 0) (–2, 2) (0, –1) so we look at the x values, like for the first one, x =-1 so it is <0 so we apply the first equation and it does not make sense because y=-x and x=-1 but y =2 not -x=1 we do that for each pair, checking the x value, finding out what equation to look at, and then see if it makes sense. with (-2,2) since x=-2<0 we look at the first equation y=-x and sure enough y=2=-x=-(-2)